2 masses are suspended by a cord that passes of rape Ollie with negligible mass. The court also has negligible mass. One the masses M1 has a mass of 7.0 kg and the other mass and to has a massive 3.0 kg. The pulley turns on a shaft through the center of the pulley, which supports the pulley and all the masses. The vertical force of the shaft on the pulley is

a. 49N upward
b. 49N downward
c. 82N upward
d. 82N downward
e. 98N upward
Answer #1

Page Ans: The situation descsibed in problem is shown below in figuse - Light Pulley Light string | Cord M2 Mi With My = 7 kgPage Fosces acting (2) I Continued com My area Mag Vestically downwards vestically upwards due to string. As the block M₂ movContinued MIT - M, Mag - (M. Mag - M₂T) = M, M₂ a M, M, a MIT - M.Mag-MM29 +MI = 0 (M +M) T = 2MM24 Putting this T= 2 M, M 2Problem is solved using basics of forces and first and second law of Newton.